OpenStudy (anonymous):

Solve the following system. x − y = 10 x + y = 8 a. (-9, 1) b. (9, -1) c. (9, 1) d. (-1, 9)

OpenStudy (ganpat):

x − y = 10 x + y = 8 solve them , adding them 2x = 18, x= 9 and y = 8 - 9 = -1 so (9, -1).. option b

OpenStudy (anonymous):

Thank you